A 38.0-kg child swings in a swing supported by two chains, each 2.90 m long. The tension in each chain at the lowest point is 436 N.



(a) Find the child's speed at the lowest point.

__ m/s





(b) Find the force exerted by the seat on the child at the lowest point. (Ignore the mass of the seat.)

___ N (upward)
1
Expert's answer
2021-10-06T15:53:14-0400

m=38.0 kg

l=2.90 m

T=436 N

(a) Child perform circular motion, so centripetal force acts on child

Force body diagram:



Centripetal force acts along the length of chain towards the support.

According to Newton's second law,

"F_{net} = centripetal \; force = \frac{mv^2}{r} \\

T+T-mg=\frac{mv^2}{r} \\

2T-mg = \frac{mv^2}{r} \\

v^2 = \frac{r(2T-mg)}{m}"

r = radius of the circular path = length of chain = 2.90 m

"v^2 = \frac{2.90(2 \times 436-38 \times 9.81}{38} \\

v^2= 38.098 \\

v=6.17 \;m/s"

(b)

"N’ -mg = \frac{mv^2}{r} \\

2T-mg = \frac{mv^2}{r} \\

2T=N’ \\

N’ = 2 \times 436 = 872 \;N"

The force exerted by the seat on the child at the lowest point is 872 N.